OpenStudy (anonymous):

MEDALS PLEASE HELP!! Please subtract and simplify 4y/y^2-6y+8 - 16/y^2-6y+8

OpenStudy (anonymous):

\[\frac{ 4_{y} }{ y ^{2}-6_{y}+8 } - \frac{ 16 }{ y ^{2}-6_{y}+8 }\] This right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Since the denominators are exactly the same, wouldn't the equation just be \[\frac{ 4_{y} -16 }{ y ^{2}-6_{y} + 8}\]

OpenStudy (anonymous):

It told me it has to be simplified down

OpenStudy (anonymous):

rheaanderson341 is correct ^

OpenStudy (anonymous):

So the correct answer is 4y-16/y^2-6y+8 ???

OpenStudy (anonymous):

No, it still can be simplified. You can simplify the \[4_{y} and 6_{y}\] along with the 16 and 8.

OpenStudy (anonymous):

\[4\div(y-2) simplified\]

OpenStudy (anonymous):